OpenStudy (anonymous):

how to explain 240:(10*4) = 240:10:4? The question is how to get from the left side to the right side? (The answer because they are both equal to the same value is not accepted)

OpenStudy (fifciol):

to divide a and b is to multiply a by the inverse of b so: \[\frac{ a }{ b }=a*\frac{ 1 }{ b }\] \[\frac{ a }{ \frac{ b }{ c } }=\frac{ a }{ b }*\frac{ 1 }{ c }\] \[\frac{ a }{ \frac{ b }{ \frac{ c }{ d } } }=\frac{ a }{ \frac{ b }{ c } }*\frac{ 1 }{ d }=\frac{ a }{ \frac{ bd }{ c } }=\frac{ a }{ bd}*\frac{ 1 }{ c }=\frac{ a }{ bcd }\] and so on

OpenStudy (fifciol):

so\[\frac{ 240 }{ \frac{ 10 }{ 4 } }=\frac{ 240 }{ 10 }*\frac{ 1 }{ 4 }=\frac{ 240 }{ 10*4 }\]